Operations Report – March 06th

  

Crew 223 Operations Report 06/03/2020

SOL: 5

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:
Off at 4:30am
On at 9:10pm SOC was 84%

Friday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 75%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149721.59 units

Water (static tank) – 307 gallons

Static to Loft Pump used – yes

Water in GreenHab – 153.7 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: yes
Hours: 149.7
Beginning charge: 100%
Ending charge: 98%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 133.7
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.5
Beginning charge: 100%
Ending charge: 80%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Honda started during the EVA this morning, and we had no
problem with it. I started the 4 ATVs during the engineer EVA for
battery check.
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? The HabCar is parked at the MDRS Campus.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA.
No problem.

Summary of Hab operations: One of the gaz button of the gaz stove is
broken (the one for the upper left).

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Florian built the solar hoven.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report